# Prompt Attacks Reveal Superficial Knowledge Removal in Unlearning Methods

### v1

@@ Appendix B Evaluation of the Rephrased WMDP-bio Datasets
To evaluate model robustness and unlearning performance, we apply a variety of input perturbation strategies to reformat multiple-choice questions in the WMDP-bio dataset. In this section, we present the results for the techniques envolving filler, rephrasing, text replacement, or translation (see Appendix A).
@@
Data Prompt Accuracy Accuracy (answered) %-Answered Accuracy using logits Accuracy using logits (right format) Accuracy using logits (wrong format)
tinyMMLU 0.5900 0.6211 0.9500 0.6200 0.6211 0.6000
WMDP 0.1430 0.3872 0.3692 0.3024 0.3894 0.2516
WMDP latin_filler_text 0.1925 0.4063 0.4737 0.3339 0.4046 0.2701
WMDP english_filler_text 0.1493 0.4158 0.3590 0.3229 0.4158 0.2708
WMDP hindi_filler_text 0.1540 0.3920 0.3928 0.3071 0.3900 0.2536
WMDP rephrased_conversation 0.0935 0.4118 0.2270 0.2891 0.4118 0.2530
WMDP rephrased_poem 0.1308 0.4368 0.2994 0.3113 0.4395 0.2565
WMDP replaced_with_variables 0.1225 0.3636 0.3370 0.2852 0.3636 0.2453
WMDP technical_terms_removed 0.1194 0.3878 0.3079 0.2969 0.3903 0.2554
WMDP translated_arabic 0.1485 0.3600 0.4124 0.2985 0.3562 0.2580
WMDP translated_bengali 0.1414 0.3346 0.4226 0.2844 0.3364 0.2463
WMDP translated_bengali 0.1335 0.3041 0.4391 0.2742 0.3059 0.2493
WMDP translated_czech 0.1225 0.4041 0.3032 0.2907 0.4041 0.2413
WMDP translated_farsi 0.1461 0.3563 0.4101 0.3064 0.3563 0.2716
WMDP translated_french 0.1259 0.4240 0.2969 0.2969 0.4240 0.2432
WMDP translated_german 0.1170 0.3716 0.3150 0.2836 0.3716 0.2431
WMDP translated_hindi 0.1587 0.3137 0.5059 0.2804 0.3152 0.2448
WMDP translated_hindi 0.1760 0.3409 0.5161 0.3009 0.3425 0.2565
WMDP translated_korean 0.1045 0.3376 0.3095 0.2828 0.3350 0.2594
WMDP translated_turkish 0.1155 0.3703 0.3119 0.2844 0.3627 0.2489
WMDP translated_vietnamese 0.1296 0.3689 0.3512 0.2992 0.3689 0.2615
@@
Data Prompt Accuracy Accuracy (answered) %-Answered Accuracy using logits Accuracy using logits (right format) Accuracy using logits (wrong format)
tinyMMLU 0.6100 0.6289 0.9700 0.6400 0.6289 1.0000
WMDP 0.1830 0.3046 0.9403 0.3018 0.3037 0.2714
WMDP latin_filler_text 0.3920 0.3986 0.9835 0.3998 0.4010 0.3333
WMDP english_filler_text 0.3912 0.3956 0.9890 0.3975 0.3979 0.3571
WMDP hindi_filler_text 0.5727 0.5800 0.9874 0.5711 0.5760 0.1875
WMDP rephrased_conversation 0.2718 0.3168 0.8578 0.3103 0.3159 0.2762
WMDP rephrased_poem 0.3231 0.3283 0.9842 0.3255 0.3267 0.2500
WMDP replaced_with_variables 0.2097 0.3152 0.6654 0.2883 0.3152 0.2347
WMDP technical_terms_removed 0.1987 0.2987 0.6654 0.2820 0.2928 0.2606
WMDP translated_arabic 0.2922 0.3032 0.9639 0.2985 0.3024 0.1957
WMDP translated_bengali 0.2844 0.2903 0.9796 0.2883 0.2879 0.3077
WMDP translated_czech 0.2694 0.2922 0.9222 0.2883 0.2922 0.2424
WMDP translated_farsi 0.3221 0.3285 0.9804 0.3284 0.3285 0.3200
WMDP translated_french 0.2692 0.2946 0.9137 0.2961 0.2955 0.3028
WMDP translated_german 0.2364 0.2954 0.8005 0.2899 0.2964 0.2638
WMDP translated_hindi 0.2828 0.2887 0.9796 0.2930 0.2903 0.4231
WMDP translated_korean 0.2608 0.3063 0.8515 0.3079 0.3072 0.3122
WMDP translated_turkish 0.2742 0.2993 0.9159 0.3064 0.3002 0.3738
WMDP translated_vietnamese 0.2864 0.3046 0.9403 0.3018 0.3037 0.2714

### v2

@@ Appendix B Evaluation of the Rephrased WMDP-bio Datasets
To evaluate model robustness and unlearning performance, we apply a variety of input perturbation strategies to reformat multiple-choice questions in the WMDP-bio dataset. In this section, we present the results for the techniques involving filler, rephrasing, text replacement, or translation (see Appendix˜A).
@@
Data Prompt acc acc (answered) %-acc acc (logits) acc (logits) (right format) acc (logits) (wrong format) tinyMMLU - 0.5900 0.6211 0.9500 0.6200 0.6211 0.6000 WMDP - 0.1430 0.3872 0.3692 0.3024 0.3894 0.2516 WMDP latin_filler_text 0.1925 0.4063 0.4737 0.3339 0.4046 0.2701 WMDP english_filler_text 0.1493 0.4158 0.3590 0.3229 0.4158 0.2708 WMDP hindi_filler_text 0.1540 0.3920 0.3928 0.3071 0.3900 0.2536 WMDP rephrased_conversation 0.0935 0.4118 0.2270 0.2891 0.4118 0.2530 WMDP rephrased_poem 0.1308 0.4368 0.2994 0.3113 0.4395 0.2565 WMDP replaced_with_variables 0.1225 0.3636 0.3370 0.2852 0.3636 0.2453 WMDP technical_terms_removed 0.1194 0.3878 0.3079 0.2969 0.3903 0.2554 WMDP translated_arabic 0.1485 0.3600 0.4124 0.2985 0.3562 0.2580 WMDP translated_bengali 0.1414 0.3346 0.4226 0.2844 0.3364 0.2463 WMDP translated_bengali 0.1335 0.3041 0.4391 0.2742 0.3059 0.2493 WMDP translated_czech 0.1225 0.4041 0.3032 0.2907 0.4041 0.2413 WMDP translated_farsi 0.1461 0.3563 0.4101 0.3064 0.3563 0.2716 WMDP translated_french 0.1259 0.4240 0.2969 0.2969 0.4240 0.2432 WMDP translated_german 0.1170 0.3716 0.3150 0.2836 0.3716 0.2431 WMDP translated_hindi 0.1587 0.3137 0.5059 0.2804 0.3152 0.2448 WMDP translated_hindi 0.1760 0.3409 0.5161 0.3009 0.3425 0.2565 WMDP translated_korean 0.1045 0.3376 0.3095 0.2828 0.3350 0.2594 WMDP translated_turkish 0.1155 0.3703 0.3119 0.2844 0.3627 0.2489 WMDP translated_vietnamese 0.1296 0.3689 0.3512 0.2992 0.3689 0.2615
@@
Data Prompt acc acc (answered) %-acc acc (logits) acc (logits) (right format) acc (logits) (wrong format) tinyMMLU - 0.6100 0.6289 0.9700 0.6400 0.6289 1.0000 WMDP - 0.1830 0.3046 0.9403 0.3018 0.3037 0.2714 WMDP latin_filler_text 0.3920 0.3986 0.9835 0.3998 0.4010 0.3333 WMDP english_filler_text 0.3912 0.3956 0.9890 0.3975 0.3979 0.3571 WMDP hindi_filler_text 0.5727 0.5800 0.9874 0.5711 0.5760 0.1875 WMDP rephrased_conversation 0.2718 0.3168 0.8578 0.3103 0.3159 0.2762 WMDP rephrased_poem 0.3231 0.3283 0.9842 0.3255 0.3267 0.2500 WMDP replaced_with_variables 0.2097 0.3152 0.6654 0.2883 0.3152 0.2347 WMDP technical_terms_removed 0.1987 0.2987 0.6654 0.2820 0.2928 0.2606 WMDP translated_arabic 0.2922 0.3032 0.9639 0.2985 0.3024 0.1957 WMDP translated_bengali 0.2844 0.2903 0.9796 0.2883 0.2879 0.3077 WMDP translated_czech 0.2694 0.2922 0.9222 0.2883 0.2922 0.2424 WMDP translated_farsi 0.3221 0.3285 0.9804 0.3284 0.3285 0.3200 WMDP translated_french 0.2692 0.2946 0.9137 0.2961 0.2955 0.3028 WMDP translated_german 0.2364 0.2954 0.8005 0.2899 0.2964 0.2638 WMDP translated_hindi 0.2828 0.2887 0.9796 0.2930 0.2903 0.4231 WMDP translated_korean 0.2608 0.3063 0.8515 0.3079 0.3072 0.3122 WMDP translated_turkish 0.2742 0.2993 0.9159 0.3064 0.3002 0.3738 WMDP translated_vietnamese 0.2864 0.3046 0.9403 0.3018 0.3037 0.2714
@@
Model Task Accuracy Zephyr_RMU tinyMMLU 0.6082 Zephyr_RMU wmdp_bio 0.3071 Zephyr_RMU wmdp_cyber 0.2718 Zephyr_RMU wmdp_chem 0.4485 Zephyr_RMU rephrased_english_filler 0.3142 Zephyr_RMU rephrased_hindi_filler 0.3009 Zephyr_RMU rephrased_latin_filler 0.3417 Zephyr_RMU rephrased_conversation 0.3040 Zephyr_RMU rephrased_poem 0.3215 Zephyr_RMU rephrased_replace_with_variables 0.2820 Zephyr_RMU rephrased_technical_terms_removed_1 0.3071 Zephyr_RMU wmdp_bio_rephrased_translated_arabic 0.3087 Zephyr_RMU wmdp_bio_rephrased_translated_bengali 0.2624 Zephyr_RMU wmdp_bio_rephrased_translated_czech 0.2907 Zephyr_RMU rephrased_translated_farsi 0.3016 Zephyr_RMU wmdp_bio_rephrased_translated_french 0.3032 Zephyr_RMU rephrased_translated_german 0.2922 Zephyr_RMU wmdp_bio_rephrased_translated_hindi 0.2899 Zephyr_RMU rephrased_translated_korean 0.2946 Zephyr_RMU wmdp_bio_rephrased_translated_telugu 0.2702 Zephyr_RMU wmdp_bio_rephrased_translated_turkish 0.2859 Zephyr_RMU wmdp_bio_rephrased_translated_vietnamese 0.2720